kamryn921:

Rectangle TUVW is on a coordinate plane at T (a, b), U (a + 2, b + 2), V (a + 5, b − 1), and W (a + 3, b − 3). What is the slope of the line that is parallel to the line that contains side UV? Answers: −2 2 −1 1

Arieonna:

@questioncovebot @supie @axie they can try to maybe help you

kamryn921:

it was wrong but thank you for helping anyways!

kamryn921:

it might have been for a different problem, but its ok!

Arieonna:

it wasn't i read the hole problem on there and it said the same esact thing this said

KyledaGreat:

We can find the slope of the line that contains side UV by using the slope formula: slope = (change in y)/(change in x) The change in y from U to V is (b-1)-(b+2) = -3, and the change in x is (a+5)-(a+2) = 3. So the slope of the line containing side UV is: slope = (-3)/3 = -1 Since we are looking for the slope of a line parallel to the line containing side UV, the slope will be the same as -1. Therefore, the answer is -1.